# A Review of Large Language Models and Autonomous Agents in Chemistry

### v1

@@ 6 Challenges and opportunities
LLMs are poised to transform fields such as drug discovery, materials science, and environmental chemistry due to their ability to predict chemical properties and reactions with remarkable accuracy. Models based on architectures like BERT have demonstrated their capability to achieve state-of-the-art performance in various property prediction tasks.43, 182 Furthermore, studies by Jablonka et al. 275 and Born and Manica 489 have showcased the predictive power of LLMs by reformulating traditional regression and classification tasks as generative tasks, opening up new avenues for chemical modeling. However, as emphasized by Weng 351, maintaining the reliability of LLM outputs is essential, as inaccuracies in formatting, logical reasoning, or content can significantly impede their practical utility. Hallucination is also an intrinsic issue with LLMs.490 Though agents can deal with hallucinations to some extent by implementing sanity-checking tools, it does not make the response hallucination-proof. A possible approach to address this issue is to use a human-in-the-loop approach, where steps of human-agent interaction are added to the workflow to check if the agent is in the correct pathway to solve the request.491, 492, 493
The potential of LLMs to design novel molecules and materials was highlighted by the AI-powered robotic lab assistant, A-Lab, which synthesized 41 new materials within just 17 days.494 Nonetheless, this achievement has sparked debates about the experimental methods and the actual integration of atoms into new crystalline materials, raising questions about the authenticity of the synthesized structures.495 These controversies underline the necessity for rigorous standards and the critical role of human expertise in validating AI-generated results. Again, the integration of advanced AI tools with the oversight of seasoned chemists is crucial, suggesting that a hybrid approach could significantly enhance both the innovation and integrity of materials science research.
In parallel, we have seen how LLM-based agents are increasingly capable of automating routine tasks in chemical research, which traditionally consume significant time and resources. These models excel in real-time data processing, managing vast datasets, and even conducting comprehensive literature reviews with minimal human intervention. Advances in AI technology now allow agents not only to perform predefined tasks but also to adapt and develop new tools for automating additional processes. For instance, tasks such as data analysis, literature review, and elements of experimental design are now being automated.286, 496, 497, 430, 435, 433, 45, 46 This automation liberates chemists to focus on more innovative and intellectually engaging aspects of their work, and the opportunity is to expand productivity and creativity in their science.
@@ Acknowledgments
M.C.R. and A.D.W. acknowledge the U.S. Department of Energy, Grant No. DE-SC0023354, and C.J.C. gratefully acknowledges the Jane King Harris Endowed Professorship at Rochester Institute of Technology for the support provided for this publication. We are grateful for feedback from early drafts of this review from the following colleagues: Kevin Jablonka, Philippe Schwaller, Michael Pieler, and Ryan-Rhys Griffiths.

### v2

@@ 6 Challenges and opportunities
LLMs are poised to transform fields such as drug discovery, materials science, and environmental chemistry due to their ability to predict chemical properties and reactions with remarkable accuracy. Models based on architectures like BERT have demonstrated their capability to achieve state-of-the-art performance in various property prediction tasks.45, 155 Furthermore, studies by Jablonka et al. 249 and Born and Manica 145 have showcased the predictive power of LLMs by reformulating traditional regression and classification tasks as generative tasks, opening up new avenues for chemical modeling. However, as emphasized by Weng 375, maintaining the reliability of LLM outputs is essential, as inaccuracies in formatting, logical reasoning, or content can significantly impede their practical utility. Hallucination is also an intrinsic issue with LLMs.519 Though agents can deal with hallucinations to some extent by implementing sanity-checking tools, it does not make the response hallucination-proof. A possible approach to address this issue is to use a human-in-the-loop approach, where steps of human-agent interaction are added to the workflow to check if the agent is in the correct pathway to solve the request.520, 521, 522
The potential of LLMs to design novel molecules and materials was highlighted by the AI-powered robotic lab assistant, A-Lab, which synthesized 41 new materials within just 17 days.523 Nonetheless, this achievement has sparked debates about the experimental methods and the actual integration of atoms into new crystalline materials, raising questions about the authenticity of the synthesized structures.524 These controversies underline the necessity for rigorous standards and the critical role of human expertise in validating AI-generated results. Again, the integration of advanced AI tools with the oversight of seasoned chemists is crucial, suggesting that a hybrid approach could significantly enhance both the innovation and integrity of materials science research.
In parallel, we have seen how LLM-based agents are increasingly capable of automating routine tasks in chemical research, which traditionally consume significant time and resources. These models excel in real-time data processing, managing vast datasets, and even conducting comprehensive literature reviews with minimal human intervention. Advances in AI technology now allow agents not only to perform predefined tasks but also to adapt and develop new tools for automating additional processes. For instance, tasks such as data analysis, literature review, and elements of experimental design are now being automated.309, 525, 526, 456, 461, 459, 47, 48 This automation liberates chemists to focus on more innovative and intellectually engaging aspects of their work, and the opportunity is to expand productivity and creativity in their science.
@@ Acknowledgments
M.C.R. and A.D.W. acknowledge the U.S. Department of Energy, Grant No. DE-SC0023354, and C.J.C. gratefully acknowledges the Jane King Harris Endowed Professorship at Rochester Institute of Technology for the support provided for this publication. We are grateful for feedback from early drafts of this review from the following colleagues: Kevin Jablonka, Philippe Schwaller, Michael Pieler, Ryan-Rhys Griffiths, Geemi Wellawatte, and Mario Krenn.